OpenStudy (anonymous):

what is cot(pi/6) in radians?

OpenStudy (ghazi):

it is equal to 1/ tan pi/6= 1/1/ sqrt 3= sqrt3\[\sqrt 3\]

OpenStudy (anonymous):

it's just square root of 3?

OpenStudy (zzr0ck3r):

im confused, this is not equal to something in radians, the radians are the input. The output is not radians. pi/6 is radians

OpenStudy (anonymous):

well that is the answer not in decimal form?